OpenStudy (anonymous):

14(1.042)^n=4.5(1.092)^n what is n?

OpenStudy (anonymous):

sorry. its 1.094

OpenStudy (anonymous):

14(1.042)^n=4.5(1.094)^n

OpenStudy (kropot72):

For your first step in solving take logs of both sides. Can you do that?

OpenStudy (anonymous):

14/4.5=(1.094/1.042)^n and then if we do log it will be log (14/4.5)=n log(1.094/1.042)

OpenStudy (anonymous):

n=23.30613587 wright?

OpenStudy (kropot72):

Now you can divide both sides of your equation by log(1.094/1.042) to find the value of n.

OpenStudy (kropot72):

Yes, your result is correct. Well done!